# GAMA: A Neural Neighborhood Search Method with Graph-aware Multi-modal Attention for Vehicle Routing Problem

### v1

@@ 2 Related Work
Recently, learning-based approaches have emerged as a promising alternative by enabling data-driven, adaptive decision-making. Within the Learning-to-Optimize (L2O) paradigm, three main sub-fields have gained prominence: learning-to-construct (L2C), learning-to-predict (L2P) solvers, and learning-to-improve (L2I). Related literature can be found in the appendix A.1.
@@ 4.1 Setup
where $d(I_{m})$ represents the total distance of the solution for the $m$-th test instance.
@@ 4.3 Results and Discussions
The result is average total distance over 500 test instances, which is calculated as Eq.equation 10, and the value is the smaller the better. Table 1 presents the performance comparison of all algorithms on CVRP instances of sizes 20, 50, and 100. We report the best objective value, average objective value over 30 independent runs, and total wall-clock time for each method.
@@
LKH3 6.0867 2h 10.3879 7h 15.6752 12h
HGS 6.0807 6.0812 45m 10.3515 10.3548 2h 15.6590 15.6994 4.5h
VNS 6.0827 6.0844 1.5h 10.4140 10.4199 13.5h 15.8843 15.8940 2.6d
POMO (gr.) 6.1111 6.1768 1s 10.5062 10.5702 4s 15.7936 15.8451 20s
POMO (A=8) 6.0904 6.1413 5s 10.4472 10.4930 30s 15.7337 15.7863 2m
LEHD (gr.) 6.3823 6.3946 3s 10.7617 10.7785 13s 17.3004 17.3188 40s
LEHD (RRC=1000) 6.0904 6.0915 30m 10.4771 10.4856 1.1h 15.8419 15.8514 2.9h
ReLD (gr.) 6.1309 6.1401 1s 10.4547 10.4676 3s 15.7558 15.7728 15s
ReLD (A=8) 6.1001 6.1041 5s 10.3877 10.3958 30s 15.6493 15.6593 2m
DACT (T=5k) 6.0811 6.0817 1m 10.3966 10.4038 3m 15.7906 15.8030 8m
DACT (T=10k) 6.0808 6.0813 2m 10.3662 10.3735 6m 15.7321 15.7410 16m
DACT (T=20k) 6.0808 6.0811 4m 10.3513 10.3542 9m 15.6853 15.6925 33m
L2I (T=5k) 6.0831 6.0864 4h 10.4012 10.4310 9.1h 15.8003 15.8914 1.6d
L2I (T=10k) 6.0815 6.0835 8h 10.3803 10.4006 18.3h 15.7207 15.8008 3.2d
L2I (T=20k) 6.0810 6.0820 16.5h 10.3607 10.3787 1.52d 15.6663 15.7334 6.5d
GIRE (T=5k) 6.1304 6.1394 20m 10.3821 10.4105 59m 15.6980 15.7453 2.3h
GIRE (T=10k) 6.1301 6.1352 41m 10.3754 10.3848 2h 15.6562 15.7023 4.6h
GIRE (D2A=5,T=20k) 6.1301 6.1322 6.8h 10.3671 10.3714 19.7h 15.5863 15.6315 1.9d
GAMA (T=5k) 6.0823 6.0836 4.5h 10.3966 10.4057 10h 15.7339 15.7389 1.6d
GAMA (T=10k) 6.0810 6.0818 9h 10.3711 10.3742 19h 15.6512 15.7054 3.3d
GAMA (T=20k) 6.0806 6.0810 19h 10.3512 10.3533 1.6d 15.6178 15.6510 6.6d

### v2

@@ 2 Related Work
Recently, learning-based approaches have emerged as a promising alternative by enabling data-driven, adaptive decision-making. Within the Learning-to-Optimize (L2O) paradigm, three main sub-fields have gained prominence: learning-to-construct (L2C) luo2025boosting, learning-to-predict (L2P) solvers, and learning-to-improve (L2I) ma2022efficient; hottung2025neural; ma2023learning; ouyang2025learning. Related literature can be found in the appendix A.1.
@@ 4.1 Setup
where $d(I_{m})$ represents the total distance of the solution for the $m$-th test instance. The training time of our GAMA varies with problem sizes, i.e., around 1 day for $N=20$, 3 days for $N=50$, and 7 days for $N=100$.
@@ 4.3 Results and Discussions
The result is average total distance over 500 test instances, which is calculated as Eq.equation 10, and the value is the smaller the better. Table 1 presents the performance comparison of all algorithms on CVRP instances of sizes 20, 50, and 100. We report the best objective value, average objective value over 30 independent runs, and run one instance average cpu time for each method.
@@
LKH3 6.0867 14s 10.3879 53s 15.6752 1.95m
HGS 6.0807 6.0812 7s 10.3515 10.3548 27s 15.6590 15.6994 59s
VNS 6.0827 6.0844 43s 10.4140 10.4199 3.2m 15.8843 15.8940 17m
POMO (gr.) 6.1111 6.1768 0.98s 10.5062 10.5702 1.5s 15.7936 15.8451 2.7s
POMO (A=8) 6.0904 6.1413 1.3s 10.4472 10.4930 4.5s 15.7337 15.7863 7s
LEHD (gr.) 6.3823 6.3946 1.5s 10.7617 10.7785 3s 17.3004 17.3188 4s
LEHD (RRC=1000) 6.0904 6.0915 35s 10.4771 10.4856 1.6m 15.8419 15.8514 4m
ReLD (gr.) 6.1309 6.1401 0.06s 10.4547 10.4676 0.1s 15.7558 15.7558 0.21s
ReLD (A=8) 6.1001 6.1041 0.09s 10.3877 10.3958 0.25s 15.6493 15.6593 0.72s
DACT (T=5k) 6.0811 6.0817 55s 10.3966 10.4038 1.8m 15.7906 15.8030 2.54m
DACT (T=10k) 6.0808 6.0813 2.1m 10.3662 10.3735 3.5m 15.7321 15.7410 9.5m
DACT (T=20k) 6.0808 6.0811 4.4m 10.3513 10.3542 11.2m 15.6853 15.6925 19.3m
L2I (T=5k) 6.0831 6.0864 27s 10.4012 10.4310 1.1m 15.8003 15.8914 4.6m
L2I (T=10k) 6.0815 6.0835 57s 10.3803 10.4006 2.19m 15.7207 15.8008 9.2m
L2I (T=20k) 6.0810 6.0820 1.9m 10.3607 10.3787 4.37m 15.6663 15.7334 18.7m
GAMA (T=5k) 6.0823 6.0836 32.5s 10.3966 10.4057 1.2m 15.7339 15.7389 4.6m
GAMA (T=10k) 6.0810 6.0818 1.1m 10.3711 10.3742 2.3m 15.6512 15.7054 9.5m
GAMA (T=20k) 6.0806 6.0810 2.3m 10.3512 10.3533 4.6m 15.6178 15.6510 19m
@@ A.4.3 Generalization on benchmark datasets
(a) Time Performance
(b) Gap Performance
Figure 5: Performance comparison between different methods on VRP instances. Left: Computation time. Right: Performance gap to baseline.
A.4.4 Performance Analysis
We also conducted a comprehensive performance evaluation of all compared methods, summarized in Figure 5. The results strongly validate the superior capability of the GAMA approach in solving VRPs.
  • Superior Solution Quality: As depicted in Figure 5 (b), GAMA almostly achieved the best solution quality (smallest gap to the baseline) across all tested instances. Notably, on the small CVRP20, CVRP50, and CVRP100 instances, where LKH3 serves as the baseline, GAMA achieved an average gap of $\mathbf{-0.194\%}$, making it the only method to outperform the LKH3 baseline on average and on every size dataset. This confirms GAMA’s high search efficiency, even against strong exact heuristics on small-to-medium problems. Furthermore, on the larger CVRPLIB instances (benchmarked against BKS), GAMA’s average gap is significantly lower than all comparative methods, demonstrating excellent robustness and generalization on large-scale VRP instances.
  • Balanced Performance and Efficiency: As shown in Figure 5 (a), despite GAMA’s complex search strategy aimed at optimizing solution quality, its computational overhead remains acceptable. GAMA’s average running time is on par with other L2I methods (such as DACT, L2I). It is reasonable and acceptable to spend some time to improve quality.
CVRP1000 CVRP2000
Best Avg Time Best Avg Time
POMO 93.8710 143.5383 30s 396.3919 486.5468 80s
POMO(A=8) 59.1911 87.6515 3.2m 169.8233 285.5397 43m
LEHD 48.5214 46.2542 32s 158.65 149.54 8m
LEHD(RRC=1000) 37.6219 38.5661 2.3h 117.52 122.4589 5.9h
ReLD 38.7953 39.0602 8s 63.3600 63.6378 33s
ReLD(A=8) 38.4513 38.6234 44s 61.9754 62.3018 2.92m
DACT(T=5k) 45.7543 46.3442 4h 70.6339 77.4895 10h
DACT(T=10k) 44.7730 45.7689 8.1h 70.6339 72.5638 20.5h
DACT(T=20k) 44.1439 45.0561 16h 68.1593 69.7607 41h
L2I(T=5k) 47.0318 68.431 18m 74.3857 153.071 1.5h
L2I(T=10k) 45.1571 62.1688 36m 70.3181 132.6643 3h
L2I(T=20k) 45.1571 57.872 1.2h 70.3181 116.3726 6h
GAMA(T=5k) 36.9435 37.2608 22.5m 58.5433 60.0774 1.8h
GAMA(T=10k) 36.7561 37.0043 44.8m 57.9618 58.7187 3.6h
GAMA(T=20k) 36.7561 36.7768 1.5h 57.9618 58.0593 7.2h
A.4.5 Evaluation on Large-scale VRP
We further evaluated the performance of our proposed GAMA method on large-scale VRP instances (CVRP1000, Capacity=250; CVRP2000, Capacity=300), following the suggestions by luo2025boosting. The evaluation results are presented in Table 8. The analysis demonstrates that GAMA maintains the best solution quality, even when faced with large-scale problems. Across both CVRP1000 and CVRP2000 datasets, GAMA(T=20k) consistently achieved the Best and Avg results among all comparative methods. For instance, on CVRP1000, GAMA(T=20k)’s average tour length was 36.7768, significantly outperforming the next best ReLD; similarly, on CVRP2000, the average solution length substantially surpassed other L2Opt methods’ performance. This strongly validates GAMA’s generalization ability and robustness for tackling large-scale VRP instances.